OpenStudy (sissyedgar):

What is the value of w? 5(9 − 3) = w • 9 – 5 • 3 I thinkits... 5

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

where did you go?

OpenStudy (solomonzelman):

I hope you see this when you come back. Just giving an abstract example: \(\large\color{black}{ \displaystyle a(b-c)=xb-ac }\) \(\large\color{black}{ \displaystyle ab-ac=xb-ac }\) \(\large\color{black}{ \displaystyle ab=xb }\) \(\large\color{blue}{ \displaystyle a=x }\)